OpenStudy (anonymous):

2logx12-4logx3=2 x is base solve for e

OpenStudy (kainui):

Well by exponent/log rules (They are basically the same thing, if you think about it) we know: a*log(b)=log(b^a) and log(a)+log(b)=log(a*b) and similarly, -log(a)=log(1/a). So using those ideas, you can combine everything together into one logx(stuff) and put both sides as powers of x.